OpenStudy (anonymous):

[Help!!<3] David built a fishpond that is shaped like a right triangle. The length of the sides of the ponds are in ratio 8:15:17. If the perimeter is 120ft, what is the length of the shorter side?

OpenStudy (anonymous):

shortest side as fraction is 8 / 40 = 1/5 length = 1/5 * 120 = 24 ft

OpenStudy (anonymous):

Let k be a multiplication factor such that 8k is the answer for the shortest side. Solve the following for k: k (8 + 15 + 17) = 120 k = 3 8 k = 24